Bacteria X has a growth rate of 160 % per hour. Some amount of bacteria X are accidentally introduced into some fruit salad. Two hours after contamination, there were 40,000 bacteria X in the fruit salad.
Find the initial number of bacteria X introduced into the fruit salad:
________bacteria
Estimate the number of bacteria in the food 3 hours after contamination.
________bacteria

Answer by ankor@dixie-net.com(22740)

Bacteria X has a growth rate of 160 % per hour.
Some amount of bacteria X are accidentally introduced into some fruit salad.
Two hours after contamination, there were 40,000 bacteria X in the fruit salad.
Find the initial number of bacteria X introduced into the fruit salad:
let Ao = initial amt of bacteria
t = 2 hrs
1.6 = the decimal form of 160% increase
Ao*1.6^2 = 40000
using the calc
Ao*2.56 = 40000
Ao = 40000/2.56
Ao = 15625 amt of bacteria initially
:
Estimate the number of bacteria in the food 3 hours after contamination.
A = 15625*1.6^3
A = 15625*4.096
A = 64000 after 3 hrs